4. Find the square roots of the following numbers by the Prime Factorisation Method. i. 729 ii. 400 iii. 1764  

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


729 = 3×3×3×3×3×3×1
⇒ 729 = (3×3)×(3×3)×(3×3) ⇒ 729 = (3×3×3)×(3×3×3) ⇒ 729 = (3×3×3)2⇒√729=3×3×3=27

400 = 2×2×2×2×5×5×1
⇒ 400 = (2×2)×(2×2)×(5×5) ⇒400 = (2×2×5)×(2×2×5)⇒ 400 = (2×2×5)2⇒√400=2×2×5=20

1764 = 2×2×3×3×7×7
⇒ 1764 = (2×2)×(3×3)×(7×7) ⇒ 1764 = (2×3×7)×(2×3×7) ⇒ 1764 = (2×3×7)2⇒√1764=2×3×7=42

Thanks